A 16 foot ladder is leaning up against a building, 7 feet from the wall. How highs the ladder?
Answer by nerdybill(7384)

A 16 foot ladder is leaning up against a building, 7 feet from the wall. How highs the ladder?
.
Let x = how high the ladder reaches
then applying Pythagorean theorem:
x^2 + 7^2 = 16^2
x^2 + 49 = 256
x^2 = 207
x = 14.39 feet